OpenStudy (anonymous):

I would like to learn how to solve this step by step please. Given f(x)=log10x and g(x)=5x-2 what is f(x)*g(x)?

OpenStudy (anonymous):

\[f(x)\cdot g(x)=\log 10x \cdot (5x-2)\\f(x)\cdot g(x)=(5x-2)\log 10x\]

OpenStudy (anonymous):

can you please tell me how you got this?

OpenStudy (anonymous):

@konrad509

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@Godlovesme

OpenStudy (squirrels):

@Luigi0210 can help you, he is very good at math!

OpenStudy (anonymous):

ok thanks ill try and contactthem ^^

OpenStudy (anonymous):

@Angel_Bear19

OpenStudy (anonymous):

@brucebaner

OpenStudy (anonymous):

@CarlyFrank07

OpenStudy (anonymous):

@dmndlife24

OpenStudy (anonymous):

@ejecktedshotgun

OpenStudy (anonymous):

@grrrxxiii

OpenStudy (anonymous):

@teleah123

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

f(x)*g(x) means multiply the two functions. So you just multiply what each function equals. log(10x) times (5x-2) equals (5x-2)log(10x). There are more ways to write this, but that's basically all you're doing.

OpenStudy (anonymous):

thank you so much i understand completely now

OpenStudy (anonymous):

@jwhite12

OpenStudy (anonymous):

one question, is (5x-2)log(10x). the same as log(10x)^5x-2 ?

OpenStudy (anonymous):

If you mean \[\log(10x)^{5x-2}\] then yes.